OpenStudy (anonymous):

A polynomial function can be written as (x − 2)(x − 3)(x + 5). What are the x-intercepts of the graph of this function?

undeadknight26 (undeadknight26):

When graphed:

undeadknight26 (undeadknight26):

So what are the x- intercepts?

OpenStudy (shamim):

U hv equation y=(x-2)(x-3)(x+5)

OpenStudy (usukidoll):

what about solving for x in each of those 3 cases?

OpenStudy (shamim):

Put y=0 in ur given equation x=?

OpenStudy (shamim):

X axis means y=0

OpenStudy (shamim):

U know either x-2=0 Or x-3=0 Or x+5=0

OpenStudy (shamim):

U will get 3 values of x=?

OpenStudy (usukidoll):

solve for x for x -2=0 , x-3=0, and x+5=0

OpenStudy (shamim):

If x-2=0 x=?

OpenStudy (shamim):

Or x-3=0 x=?

OpenStudy (shamim):

Or x+5=0 x=?

OpenStudy (arindameducationusc):

2,3,-5 I guess

OpenStudy (arindameducationusc):

The steps of @UsukiDoll @shamim are very much right...

OpenStudy (usukidoll):

^_^

OpenStudy (anonymous):

There's only 1 x-intercept. x=30.

OpenStudy (anonymous):

Oops, y-intercept.

OpenStudy (anonymous):

x intercept is (-5,0), (2,0), (3,0)